OpenStudy (anonymous):

write a equation of a line through points (5, 5) and (1 ,0) in point slope form. A) y= 1/4 (x - 5) B) y-1=5/4 (x - 1) C) y-5= 1/4 (x - 1) D) y=5/4 (x - 1)

OpenStudy (anonymous):

hiii :D

OpenStudy (anonymous):

find the slope: m=y2-y1/x2-x1 and plug into point slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

still confused explain more